OpenStudy (anonymous):

At last years company picnic, 25 pounds of ground beef was need to make 100 hamburgers. How many pounds of ground beef will be needed this year if 120 hamburgers are to be made? A. 30 Ib B. 45 Ib C. 50 Ib D. 55 Ib

OpenStudy (anonymous):

@aminamoazzam @aryandecoolest

OpenStudy (anonymous):

@Thesmarterone

OpenStudy (anonymous):

\[(25*120)/100\]

OpenStudy (anonymous):

Ok I did 25*120 and got 3000 so what would I do with the 100

OpenStudy (aminamoazzam):

25 pounds for 100 burger x pounds for 120 burger so 120*25 = 100x 3000/100 = 30 Ans = A

OpenStudy (anonymous):

yep!! right !! 3000/100=30

OpenStudy (anonymous):

Oh ok thank y'all for they help

OpenStudy (anonymous):

pleasure!!! ND ANYTIME